OpenStudy (anonymous):

The difference between a number and eleven is three more than half the number is???

OpenStudy (kropot72):

Let the number be x. Then to find x you need to solve the following equation: \[x-11+3=\frac{x}{2}\]

OpenStudy (anonymous):

so it would look like this? x - 11 = x/2 + 3

OpenStudy (kropot72):

Not really. x - 11 + 3 = x/2 Then subtracting 3 from both sides gets x - 11 = x/2 - 3